OpenStudy (anonymous):

Find the arc length of the curve x = 8et cos 6t, y = 8et sin 6t where 0 ≤ t ≤ π/2.

OpenStudy (anonymous):

i got my dx/dt to be \[-48e ^{t}\sin(6t) +8e ^{t}\cos(6t)\]

OpenStudy (anonymous):

and i got my dy/dt to be \[48e ^{t}\cos(6t) +8e ^{t}\sin(6t)\]

OpenStudy (anonymous):

-48etsin6t+48etcos6t

OpenStudy (anonymous):

what's that for?

OpenStudy (anonymous):

what are you getting for your derivatives or your g'(x) and f'(x)?

OpenStudy (anonymous):

derivative of x -48etsin6t

OpenStudy (anonymous):

derivative of y 48etcos6t

OpenStudy (anonymous):

what about the product rule?

OpenStudy (anonymous):

:D sorry

OpenStudy (anonymous):

lol np

OpenStudy (anonymous):